OpenStudy (anonymous):

Use trigonometric identities to simplify the trigonometric expression: (cscx+cotx)(cscx-cotx)

OpenStudy (anonymous):

im assuming it starts out like this... csc^2x - cot^2x

jimthompson5910 (jim_thompson5910):

hint: (a+b)(a-b) = a^2 - b^2

jimthompson5910 (jim_thompson5910):

good start, what's next?

OpenStudy (anonymous):

yeah thats where im stuck=/

jimthompson5910 (jim_thompson5910):

well one pythagorean identity is 1 + cot^2 = csc^2 and it's derived from the fact that sin^2 + cos^2 = 1

jimthompson5910 (jim_thompson5910):

how can you manipulate that to get csc^2 - cot^2 ?

jimthompson5910 (jim_thompson5910):

btw I found this (along with others) here http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

if you move the cot^2 to the other side by subtraction..

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so 1 + cot^2 = csc^2 1 + cot^2 - cot^2 = csc^2 - cot^2 1 = csc^2 - cot^2

OpenStudy (anonymous):

oh, so the answer is 1?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

wow, that was easy! thanks again!!!

jimthompson5910 (jim_thompson5910):

a very fancy way of saying 1 since it's always true for all of the defined values of x

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

alright, got it!

OpenStudy (anonymous):

oh and thanks for that link!!

jimthompson5910 (jim_thompson5910):

sure thing, that is definitely very handy (since you aren't expected to memorize them all, at least I hope)

OpenStudy (anonymous):

hah, yeah no way

jimthompson5910 (jim_thompson5910):

you never know lol, but let's hope your teacher is understanding

OpenStudy (anonymous):

god willing!